# Rheology of suspensions of non-Brownian soft spheres across the jamming and viscous-to-inertial transitions

### v2

@@ Abstract
The rheology of suspensions of non-Brownian soft spheres is studied across jamming but also across the viscous and inertial regimes using a custom pressure- and volume-imposed rheometer. The study shows that the granular rheology found for suspensions of hard spheres can be extended to a soft granular rheology (SGR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values and using the addition of the viscous and inertial stress scales. This SGR encompasses rheological behaviors on both sides of the jamming transition, resulting in an approximate collapse of the rheological data into two branches when scaled with the distance to jamming, as observed for soft colloids. This research suggests that suspensions of soft particles across flow regimes can be described by a unified SGR framework around the jamming transition.
@@
The objective of this paper is to examine the rheology of soft spherical particles across jamming but also across the viscous and inertial regimes. To this end, we use large hydrogel spheres suspended in fluids of variable viscosity and a custom pressure- and volume-imposed rheometer. We show that this soft particulate system can be described by a soft granular rheology (SGR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values [10] and using the superposition of the viscous and inertial stress scales [13, 14, 15]. We also demonstrate that the SGR is capable of capturing the rheological behaviors on both sides of the jamming transition, resulting in an approximate collapse of the rheological data into two branches when scaled with the distance to jamming, as found for soft colloids [4], but now using stress additivity across the flow regimes.
Figure 1: (a) $\phi_{c}$ versus $P$ for the case $\eta_{f}=56$ mPa s and $h=6d$ at $v=10\ \mu$m/s and $=100\ \mu$m/s [inset: blowup of the zero pressure zone defining $\phi_{c}(0)$] during the second cycle of uniaxial compression-decompression, (b) $\phi_{c}(0)$ (inferred from decompression curves) versus $\eta_{f}$, and (c) $\phi_{c}/\phi_{c}(0)-1$ versus $P/E$ for $h=6d$ at $v=10\ \mu$m/s from decompression data
@@
In Figs. 2(a)-inset and (d), $\mu$ and $\phi$ are plotted against $J+\alpha_{\phi}I^{2}$ across the jamming transition for 5 different increasing imposed pressure shown by the increased color intensity for each suspension with varying fluid viscosity. The coefficient $\alpha_{\phi}=0.1$ corresponds to the inverse of the transitional Stokes number $St_{v\rightarrow i}=10$ found for frictional hard spheres [15]. In the same way as in this later work [15], a better collapse is obtained for $\mu$ when plotted against $J+\alpha_{\mu}I^{2}$ with a smaller $\alpha_{\mu}=0.0088$ characteristic of a larger transitional Stokes number for $\mu$, $St^{\mu}_{v\rightarrow i}\approx 114$, see the comparison of the main graph and the inset of Fig. 2(a). In Fig. 2(e) and (b), the hard-sphere granular rheology [15] is extended to the SGR [10],
@@
The quasi-static values inferred from the fits, $\phi_{c}$ and $\mu_{c}$, are shown in the insets of Figs. 2(e) and (b). Their $P$-dependence can be described by power laws [10, 18],
@@
While $\mu_{c}$ shows a weak dependence and has a value $\approx 0.11$ characteristic of low-friction particles, $\phi_{c}$ presents a conspicuous increase. This latter effect is better evidenced in Fig. 3(a) where the relative value $\phi_{c}/\phi_{c}(0)-1$ is instead used to avoid the large uncertainties on the absolute value of $\phi_{c}$. The $P$-imposed data of the inset of Fig. 2(e) are complemented by data coming from $\phi$-imposed rheometry using Eq. (1). Fitting the data to Eq. (3) (black dashed curve) yields a coefficient $c_{\phi}=4.7\pm 0.2$ and an exponent $x=0.67\pm 0.02$ agreeing with Hertz law [18] as obtained at low $P/E$ in the differing experiment of Fig. 1(c) which uses uniaxial decompression. The values of $\phi_{c}(0)$ deduced from the fit are shown in Fig. 3(b) for the 5 different suspensions and $P$- and $\phi$-imposed measurements and are close to $\approx 0.64$ within ±$\pm 5\%$.
Figure 3: (a) $\phi_{c}/\phi_{c}(0)-1$ versus $P/E$ and (b) $\phi_{c}(0)$ versus $\eta_{f}$ for $P$- and $\phi$-imposed data
@@
This frictional description has its dual description shown in Fig. 2(c) and (f), in which the control parameter is $\phi/\phi_{c}(P/E)$ and the rheology is described by $\tau$ and $P$ normalized by the addition of stress scales $\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2}$. The important result shown in the insets is that the normalized $\tau$ and $P$ functions diverge as $(1-\phi/\phi_{c})^{-\frac{1}{\gamma}}$ with $\gamma=0.3$ for these low-frictional spheres instead of $(1-\phi/\phi_{c})^{-\frac{1}{0.5}}$ for frictional spheres, and thus in better agreement with the theoretical predictions for frictionless spheres [13, 26].
The SGR given by Eqs. (1-4) provides a description of the rheology of non-Brownian soft particles on both sides of the jamming transition [10], see the full calculation in the Supplemental Material 3 as only the asymptotic limits are reported below. For $\phi>\phi_{c}(0)$ in the zero shear limit, the SGR accounts for finite yield pressure and stress,
@@
$\tau_{y}/P_{y}$ $=$ $\mu_{c}(0)\left\{1+c_{\mu}\{[\phi/\phi_{c}(0)-1]/c_{\phi}\}^{\frac{y}{x}}\right\}\approx\mu_{c}(0).$ (6)
For $\phi<\phi_{c}(0)$ in the zero shear limit, the SGR rationalizes the diverging behaviors as
$P/(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})$ $\sim$ $\left\{[1-\phi/\phi_{c}(0)]/a_{\phi}\right\}^{-\frac{1}{\gamma}},$ (7)
$\tau/P$ $\sim$ $\mu_{c}(0)\left\{1+\frac{a_{\mu}}{a_{\phi}}\left[1-\phi/\phi_{c}(0)\right]\right\}.$ (8)
@@
Close to jamming, i.e. $\phi\rightarrow\phi_{c}(0)$, the SGR produces non-trivial shear-thinning behaviors with, in the present small deformation limit $c_{\phi}(P/E)^{x}\ll 1$, the approximations
$P/E$ $\sim$ $\left(a_{\phi}/c_{\phi}\right)^{\frac{1}{\gamma+x}}\left[(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})/E\right]^{\frac{\gamma}{\gamma+x}},$ (9)
$\tau/P$ $\sim$ $\mu\left(\phi_{c}(0),St,\frac{P}{E}\right),$ (10)
@@
However, the collapse of the data onto two curves (one above jamming and one below) obtained upon scaling the stresses and shear rate as power laws of the distance to jamming [6] is only approximative within this SGR framework [10], in particular for the shear stress where the asymptotic behaviors are not connected by simple power laws as seen by Eqs. (6),(8), and (10).
@@
Using the critical exponents $\Delta=1/x\,(=1.49$ within $3\%$) and $\Gamma=1/x+1/\gamma\,(=4.83$ within $30\%$) deduced from the asymptotic behaviors given by Eqs. (5-10), we plot $\tau/E|1-\phi/\phi_{c}(0)|^{\Delta}$ and $P/E|1-\phi/\phi_{c}(0)|^{\Delta}$ versus $\tilde{S}=(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})/E|1-\phi/\phi_{c}(0)|^{\Gamma}$ using the deduced $\phi_{c}(0)$ of the different suspensions (given the inset of Fig. 3) for both $P$- and $\phi$-imposed data in Fig. 4. In the insets, we show typical curves of $\tau$ and $P$ versus $\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2}$ (all normalized by $E$) across the jamming transition obtained from $\phi$-imposed measurements which are in excellent agreement with the SGR predictions (dashed curves). For $\phi>\phi_{c}(0)$, the stresses extrapolates toward a nonzero yield stress, while for $\phi<\phi_{c}(0)$, they tend towards zero at low strain rates. In the main plots of Fig. 4, a good collapse of the data onto two curves is obtained for the 5 particle-fluid mixtures. As previously mentioned, the collapse happens to be better for the $P$ data than for the $\tau$ data. The collapsed branch above $\phi_{c}(0)$ is approximately a generalized Herchel-Bulkley (H-B) law (solid curve), with yield stress given by Eqs. (5-6), and a shear-thinning exponent of $\Delta/\Gamma\,(=0.3)$, see Eqs. (9-10). The collapsed branch below $\phi_{c}(0)$ is approximately a power law with the same exponent $\approx\Delta/\Gamma$ which merges with the upper branch very close to the jamming transition (dotted line). Interestingly, this lower branch turns to a linear variation given by Eqs. (7-8) (dashed-dotted line) far from jamming as the elastic part of the spheres becomes less dominant.
@@
In conclusion, we have used a custom pressure- and volume-imposed rheometer to obtain reliable rheological data for soft hydrogels above and below the jamming transition and across the viscous and inertial flow regimes. In addition, we have characterized the mechanical properties of these suspensions using uniaxial compression-decompression experiments. The first major result is that we can generalize the granular rheology found for suspensions of hard spheres [15] to a soft granular rheology (SGR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values (weakly-dependent for the friction coefficient) and using the additivity of the viscous and inertial stress scales. The main difference is that the present particles are soft and possess a low friction, which results in their asymptotic behavior differing from that of frictional particles. The second important finding is that this SGR provides a comprehensive description of the rheological behaviors on both sides of the jamming transition and leads to an approximate collapse of the rheological data into two branches, which is analogous to the behavior observed in soft colloids [4, 5]. This is a remarkable result as it means that suspensions of soft particles across the entire range of size and flow, i.e. from the colloidal to the granular realm, can be described by the same SGR framework around the jamming transition.

### v3

@@ Abstract
The rheology of suspensions of non-Brownian soft spheres is studied across jamming but also across the viscous and inertial regimes using a custom pressure- and volume-imposed rheometer. The study shows that the granular rheology found for suspensions of hard spheres can be extended to a soft granular rheology (SGranR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values and using the addition of the viscous and inertial stress scales. This SGranR encompasses rheological behaviors on both sides of the jamming transition, resulting in an approximate collapse of the rheological data into two branches when scaled with the distance to jamming, as observed for soft colloids. This research suggests that suspensions of soft particles across flow regimes can be described by a unified SGranR framework around the jamming transition.
@@
The objective of this paper is to examine the rheology of soft spherical particles across jamming but also across the viscous and inertial regimes. To this end, we use large hydrogel spheres suspended in fluids of variable viscosity and a custom pressure- and volume-imposed rheometer. We show that this soft particulate system can be described by a soft granular rheology (SGranR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values [10] and using the superposition of the viscous and inertial stress scales [13, 14, 15, 16]. We also demonstrate that the SGranR is capable of capturing the rheological behaviors on both sides of the jamming transition, resulting in an approximate collapse of the rheological data into two branches when scaled with the distance to jamming, as found for soft colloids [4], but now using stress additivity across the flow regimes.
Figure 1: (a) $\phi_{c}$ versus $P$ for the case $\eta_{f}=56$ mPa s and $h=6d$ at $v=10\ \mu$m/s and $=100\ \mu$m/s [inset: blowup of the zero pressure zone defining $\phi_{c}(0)$] during the second cycle of uniaxial compression-decompression, (b) $\phi_{c}(0)$ (inferred from decompression curves) versus $\eta_{f}$, and (c) $\phi_{c}/\phi_{c}(0)-1$ versus $P/E$ for $h=6d$ at $v=10\ \mu$m/s from decompression data.
@@
In Figs. 2(a)-inset and (d), $\mu$ and $\phi$ are plotted against $J+\alpha_{\phi}I^{2}$ across the jamming transition for 5 different increasing imposed pressure shown by the increased color intensity for each suspension with varying fluid viscosity. The coefficient $\alpha_{\phi}=0.1$ corresponds to the inverse of the transitional Stokes number $St_{v\rightarrow i}=10$ found for frictional hard spheres [16]. In the same way as in this later work [16], a better collapse is obtained for $\mu$ when plotted against $J+\alpha_{\mu}I^{2}$ with a smaller $\alpha_{\mu}=0.0088$ characteristic of a larger transitional Stokes number for $\mu$, $St^{\mu}_{v\rightarrow i}\approx 114$, see the comparison of the main graph and the inset of Fig. 2(a). In Fig. 2(e) and (b), the hard-sphere granular rheology [16] is extended to the SGranR [10],
@@
The quasi-static values, $\phi_{c}(P/E)$ and $\mu_{c}(P/E)$, obtained by fitting each imposed-pressure curve using the aforementioned mean values of $a_{\phi},a_{\mu},$ and $\gamma$, are shown in the insets of Figs. 2(e) and (b). Their $P$-dependence can be described by power laws [10, 19],
@@
While $\mu_{c}$ shows a weak dependence and has a value $\approx 0.11$ characteristic of low-friction particles, $\phi_{c}$ presents a conspicuous increase. This latter effect is better evidenced in Fig. 3(a) where the relative value $\phi_{c}/\phi_{c}(0)-1$ is instead used to avoid the large uncertainties on the absolute value of $\phi_{c}$. The $P$-imposed data of the inset of Fig. 2(e) are complemented by data coming from $\phi$-imposed rheometry using Eq. (1). Fitting the data to Eq. (3) (black dashed curve) yields a coefficient $c_{\phi}=4.7\pm 0.2$ and an exponent $x=0.67\pm 0.02$ agreeing with Hertz law (for which $x=2/3$) [19] as obtained at low $P/E$ in the differing experiment of Fig. 1(c) which uses uniaxial decompression. The values of $\phi_{c}(0)$ deduced from the fit are shown in Fig. 3(b) for the 5 different suspensions and $P$- and $\phi$-imposed measurements and are close to $\approx 0.64$ within ±$\pm 5\%$, close to the values found in Fig. 1(b) under similar confinement.
Figure 3: (a) $\phi_{c}/\phi_{c}(0)-1$ versus $P/E$ and (b) $\phi_{c}(0)$ versus $\eta_{f}$ for $P$- and $\phi$-imposed data.
@@
The frictional description of Fig. 2(b) and (e) has its dual description shown in Fig. 2(c) and (f), in which the control parameter is $\phi/\phi_{c}(P/E)$ and the rheology is described by $\tau$ and $P$ normalized by the addition of stress scales $\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2}$, using the same $P$-imposed data. The important result shown in the insets is that the normalized $\tau$ and $P$ functions diverge as $(1-\phi/\phi_{c})^{-\frac{1}{\gamma}}$ with $\gamma=0.3$ for these low-frictional spheres instead of $(1-\phi/\phi_{c})^{-\frac{1}{0.5}}$ for frictional spheres, and thus in better agreement with the theoretical predictions for frictionless spheres [13, 27].
The SGranR given by Eqs. (1-4) provides a description of the rheology of non-Brownian soft particles on both sides of the jamming transition [10], see the full calculation in the Supplemental Material 3 as only the asymptotic limits are reported below. For $\phi>\phi_{c}(0)$ in the zero shear limit, the SGranR accounts for finite yield pressure and stress (denoted by the subscript $y$),
@@
$\tau_{y}/P_{y}$ $=$ $\mu_{c}(0)\left\{1-c_{\mu}\{[\phi/\phi_{c}(0)-1]/c_{\phi}\}^{\frac{y}{x}}\right\}\approx\mu_{c}(0).$ (6)
For $\phi<\phi_{c}(0)$ in the zero shear limit, the SGranR rationalizes the diverging behaviors as
$P/(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})$$\rightarrow$ $\left\{[1-\phi/\phi_{c}(0)]/a_{\phi}\right\}^{-\frac{1}{\gamma}},$ (7)
$\tau/P$$\rightarrow$ $\mu_{c}(0)\left\{1+\frac{a_{\mu}}{a_{\phi}}\left[1-\phi/\phi_{c}(0)\right]\right\}.$ (8)
@@
Close to jamming, i.e. $\phi\rightarrow\phi_{c}(0)$, the SGranR produces non-trivial shear-thinning behaviors with, in the present small deformation limit $c_{\phi}(P/E)^{x}\ll 1$, the approximations
$P/E$$\rightarrow$ $\left(a_{\phi}/c_{\phi}\right)^{\frac{1}{\gamma+x}}\left[(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})/E\right]^{\frac{\gamma}{\gamma+x}},$ (9)
$\tau/P$$\rightarrow$ $\mu\left(\phi_{c}(0),St,\frac{P}{E}\right),$ (10)
@@
However, the collapse of the data onto two curves (one above jamming and one below) obtained upon scaling the stresses and shear rate as power laws of the distance to jamming [6] is only approximative within this SGranR framework [10], in particular for the shear stress where the asymptotic behaviors are not connected by simple power laws as seen by Eqs. (6),(8), and (10).
@@
Using the critical exponents $\Delta=1/x\,(=1.49$ within $3\%$) and $\Gamma=1/x+1/\gamma\,(=4.83$ within $30\%$) deduced from the asymptotic behaviors given by Eqs. (5-10), we plot $\tau/E|1-\phi/\phi_{c}(0)|^{\Delta}$ and $P/E|1-\phi/\phi_{c}(0)|^{\Delta}$ versus $\tilde{S}=(\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2})/E|1-\phi/\phi_{c}(0)|^{\Gamma}$ using the deduced $\phi_{c}(0)$ of the different suspensions (given the inset of Fig. 3) for both $P$- and $\phi$-imposed data in Fig. 4. In the insets, we show typical curves of $\tau$ and $P$ versus $\eta_{f}\dot{\gamma}+\alpha_{\phi}\rho_{p}d^{2}\dot{\gamma}^{2}$ (all normalized by $E$) across the jamming transition obtained from $\phi$-imposed measurements which are in excellent agreement with the SGranR predictions (dashed curves). For $\phi>\phi_{c}(0)$, the stresses extrapolates toward a nonzero yield stress, while for $\phi<\phi_{c}(0)$, they tend towards zero at low strain rates. In the main plots of Fig. 4, a good collapse of the data onto two curves is obtained for the 5 particle-fluid mixtures. As previously mentioned, the collapse happens to be better for the $P$ data than for the $\tau$ data, in particular in the three asymptotic limits. The collapsed branch above $\phi_{c}(0)$ is approximately a generalized Herschel-Bulkley (H-B) law (solid curve), with yield stress given by Eqs. (5-6), and a shear-thinning exponent of $\Delta/\Gamma\,(=0.3)$, see Eqs. (9-10). The collapsed branch below $\phi_{c}(0)$ is approximately a power law with the same exponent $\approx\Delta/\Gamma$ which merges with the upper branch very close to the jamming transition (dotted line). Interestingly, this lower branch turns to a linear variation given by Eqs. (7-8) (dashed-dotted line) far from jamming as the elastic part of the spheres becomes less dominant.
@@
In conclusion, we have used a custom pressure- and volume-imposed rheometer to obtain reliable rheological data for soft hydrogels above and below the jamming transition and across the viscous and inertial flow regimes. In addition, we have characterized the mechanical properties of these suspensions using uniaxial compression-decompression experiments. The first major result is that we can generalize the granular rheology found for suspensions of hard spheres [16] to a soft granular rheology (SGranR) by renormalizing the critical volume fraction and friction coefficient to pressure-dependent values (weakly-dependent for the friction coefficient) and using the additivity of the viscous and inertial stress scales. The main difference is that the present particles are soft and possess a low friction, which results in their asymptotic behavior differing from that of frictional particles. The second important finding is that this SGranR provides a comprehensive description of the rheological behaviors on both sides of the jamming transition and leads to an approximate collapse of the rheological data into two branches, which is analogous to the behavior observed in soft colloids [4, 5]. This is a remarkable result as it means that suspensions of soft particles across the entire range of size and flow, i.e. from the colloidal to the granular realm, can be described by the same SGranR framework around the jamming transition.